OpenStudy (anonymous):

Solve the differential equation: (x * y' - 1) * ln(x) = 2 * y

OpenStudy (anonymous):

it's already been differentiated?

OpenStudy (anonymous):

i dont understand what you mean.

OpenStudy (anonymous):

nvm...do you know implicit differentiation?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok that makes it easier...I'll calc it in a second

OpenStudy (nowhereman):

substituting \[x := e^t\] seems to make it a lot easier

OpenStudy (anonymous):

the derivative of ln(x) is 1/x, but we have to do product rule

OpenStudy (anonymous):

sorry man not sure on this one

OpenStudy (anonymous):

and gotta get back studying for lin alg final -___-

OpenStudy (anonymous):

try www.wolframalpha.com for an answer

OpenStudy (anonymous):

oh ok thanks for your help :) @ nowhereman: if i do that i get y'*(e^t)*t=2y

OpenStudy (nowhereman):

I got \[(\frac{dy}{dt} - 1)t = 2y\]

OpenStudy (nowhereman):

assuming the function is analytic, that can be solved with power series

OpenStudy (anonymous):

would i be able to have \[(dy/dt) - (2/t)y=1\] and then do an integrating factor:\[e^(intergral of (-2/t))\]?

OpenStudy (nowhereman):

no, I don't see how that would work.

OpenStudy (anonymous):

... i dont understand what i should do then

OpenStudy (anonymous):

umm and i dont get how you got the dt part

OpenStudy (nowhereman):

I used the chain rule \[\frac{dy}{dt} = \frac{dy}{dx}\frac{dx}{dt}\]